Fermi National Laboratory
 

Accelerator Update for April 6, 2007 - April 9, 2007

 

Friday April 6

The day shift began with Tevatron (TeV) experts conducting studies, with Operations monitoring the Recycler stash (169.3E10), with the Antiproton Source (Pbar) stacking (30.7E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI and the Meson MTest experiment taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

A MI expert began a slipstacking study at 11:19 AM.

At 11:35 AM, a Linac RF station (LRF3) tripped off. It reset okay.

A Booster expert reported at 1:53 PM that he had timed in a new Booster kicker (B:MKS04), which put Booster back into its normal mode of operation with five kickers.

At 1:58 PM, a Booster RF station (BRF11) tripped off and caused the West Anode Power Supply to trip off. Experts investigated and found a failing fan. They had the fan replaced by 2:30 PM.

Operations reported at 3:57 PM that the Recycler had stashed antiprotons twice during the day shift.

Operations put the TeV through a 4x0 wet squeeze around 4 PM.

At 4:19 PM, BRF1 & 6 tripped off. They both reset.

MTest experimenters reported at 4:51 PM that they were through with beam for the day.

At 6:17 PM, Operations turned on and ran the MiniBooNE power supplies; they will be kept on until Monday morning.

Operations put the TeV into shot setup at 6:19 PM.

At 6:27 PM, the TeV suffered a sector B2 power supply dump failure. Operation put shot setup on hold and called in a power supply expert.

A power supply expert reported at 7:35 PM that the he had repaired the TeV B2 power supply. Operations conducted a successful B2 dump test at 8:01 PM.

Operations put the TeV through a dry squeeze at 8:01 PM.

Operations put the TeV into shot setup at 8:44 PM.

Operations established store 5337 at 10:57 PM with an initial luminosity of 213.04E30.

Operations reported at midnight that the Recycler had stashed twice during the evening shift.

Saturday April 7

The midnight shift began with Operations monitoring store 5337 (171.2E30) and the stash (81.3E10), with Pbar stacking (18.7E10), and with NuMI taking beam.

At 3:09 AM, a Linac Klystron RF station (KRF5) tripped off for the first of three times in seven minutes. It reset.

At 3:35 AM, BRF6 tripped off. It reset.

MTest resumed taking beam at 6:14 AM.

Operations reported 8 AM that the Recycler had stashed three times during the midnight shift.

At 8:37 AM, a Booster vacuum pump (B:IPRF17) tripped off and caused BRF7 & 8 to trip off. Everything reset.

At 12:01 PM, LRF3 tripped off for the first of seven times in six hours. It reset.

At 3:07 PM, BRF3 tripped off for the first of three times in 11 minutes. At 3:20 PM, Operations reported that the station would not stay on. Operators bypassed the station and brought BRF19 on line.

MTest reported at 6 PM that they were through with the beam for the night.

Operations terminated store 5337 at 8 PM.

Operations reported at 8:01 PM that the Recycler had stashed four times during the last 12 hours. The stack was at 37.9E10 and the stash was at 326.3E10.

Operations put the TeV into shot setup at 8:16 PM.

Operations established 5338 at 10:15 PM with an initial luminosity of 189.06E30.

At 10:41 PM, the NuMI horn tripped off. It reset.

Sunday April 8

The midnight shift began with Operations monitoring store 5338 and the stash, with Pbar stacking, and with NuMI taking beam.

At 3:38 AM, KRF3 tripped off. It reset.

At 5:08 AM, BRF6 tripped off. It reset.

MTest resumed taking beam at 6:35 AM.

Operations reported at 8:13 AM that the Recycler had stashed four times during the last 12 hours. Store 5338 was at 54.72E30, the stack was at 40.6E10, and the stash was at 162.9E10.

At 10:44 AM, the NuMI beam permit tripped off. It reset.

At 2:47 PM, LRF3 tripped off. It reset.

At 3:07 PM, BRF6 tripped off. It reset.

At 4:35 PM, KRF5 tripped off. It reset.

At 4:40 PM, a Booster vacuum pump (B:IPRF21) tripped off and caused BRF9 & 10 to trip off. Everything reset.

Operations terminated store 5338 at 5:01 PM.

Operations put the TeV into shot setup at 5:23 PM.

At 5:43 PM, a Booster vacuum pump (B:IPRF23) tripped off and caused BRF9 & 10 to trip off. Everything reset.

The MTest experimenters reported at 5:59 PM that they were done with the beam for the day.

Operations established store 5339 at 7:14 PM with an initial luminosity of 210.07E30.

At 7:28 PM, store 5339 aborted. There was no quench. Operations contacted experts.

Operations reported at 8:13 PM that the Recycler had stashed five times during the last 12 hours. The stack was at 22.2E10 and the stash was at 40.6E10.

At 8:48 PM, a MI RF station (MIRF16) tripped off and wouldn’t reset. Operators talked to an RF expert, bypassed the station, and adjusted the RF system to run without MIRF16.

TeV experts reported at 11:16 PM the TeV abort was due to a faulty sector D2 dump contacts, which they repaired.

Monday April 9

The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI taking beam.

At 12:03 AM, the I- Source chopper tripped off due to a vacuum valve closing.

At 12:13 AM, LRF3 tripped off. It reset.

At 12:21 AM, BRF6 tripped off. It reset.

At 12:47 AM, Operations reported a TeV sector A3 dump failure. It would not reset. Operations called in a power supply expert.

Operations reported at 4:15 AM that a power supply expert and operators had replaced the TeV sector A3 DC breaker. Operations began ramping the TeV.

Operations put the TeV through a dry squeeze at 4:32 AM.

Operation put the TeV through a wet squeeze at 5:35 AM.

Operations put the TeV into shot setup at 5:59 AM.

Operations established store 5342 at 7:46 AM with an initial luminosity of 147.85E30.

Operations reported at 8:14 AM that the Recycler had stashed three times during the last 12 hours.

The Plans for Monday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/09/2007    email Fermilab